OpenStudy (anonymous):

LCD of 9/12 and 3/10 . rewrite each fraction as an equivalent fraction with the LCD . Dont get this problem

OpenStudy (anonymous):

you are looking for the least common multiple of 12 and 10 it is 60 because both numbers divide 60 evenly and 60 is the smallest one

OpenStudy (anonymous):

then write \[\frac{9}{12}=\frac{5\times 9}{60}=\frac{45}{60}\] and \[\frac{3}{10}=\frac{6\times 3}{60}=\frac{18}{60}\]